# spacing=nonfrench

### v1

@@ 3.4 Dormant
Overall, to optimize the protective perturbation $\delta$ for preventing unauthorized image usage in pose-driven human image animation, we propose $\mathcal{L}_{vae}$ and $\mathcal{L}_{feature}$ for feature misextraction, $\mathcal{L}_{frame}$ for frame incoherence. We also apply several techniques to further enhance the perturbation $\delta$ during optimization, including LPIPS [79] for imperceptibility, EoT [5] for robustness and Momentum [17] for transferability.
@@
EoT. To enhance the robustness of $\delta$ against various transformations, we adopt EoT into the PGD process. EoT introduces a distribution of transformation functions to the input and optimizes the expectation of the objective function values over these transformations. To reduce time and computational costs, during each PGD step, we randomly sample one of the following transformations and apply it to the protected image $x_{p}$: Gaussian blur, JPEG compression, Gaussian noise, Random resize (resizing to a random resolution and then back to the original size), or no transformation.
@@ 4.1 Experimental Setup
Pose-driven Human Image Animation Methods. We comprehensively evaluate the effectiveness and transferability of Dormant on 8 cutting-edge and widely-used human image animation methods, including Animate Anyone [31], MagicAnimate [72], MagicPose [15], MusePose [63], Champ [86], MuseV [71], UniAnimate [69] and ControlNeXt [52].
Datasets. We conduct experiments on 4 datasets: TikTok [34], Champ [86], UBC Fashion [76] and TED Talks [62]. For the TikTok dataset, we follow the settings used in prior human image animation works [15, 68] and utilize their 10 TikTok-style videos showing different people from the web for evaluation. These videos contain between 248 and 690 frames. Specifically, we use the first frame of each video as the reference image and extract pose sequences from the remaining frames using DWPose [74] or DensePose [23] to guide the video generation. Additionally, we also sample 10 videos from the Champ training set, the UBC Fashion, and TED Talks test sets for evaluation, resulting in videos with 237-848, 303-355, and 135-259 frames, respectively. All video frames are resized to $512\times 512$, and we further conduct experiments on other image resolutions in Section 4.4.
Baseline Protections. We compare the protection performance of Dormant with 6 baseline methods defending against text/image-to-image generation (SDS [73], Glaze 2.1 [61], Mistv2 [84], PhotoGuard [60], and AntiDB [65]) and image-to-video generation (VGMShield [50]). For a fair comparison, we set the perturbation budget $\eta=16/255$ and PGD iterations $N=200$ for all protection methods [^5] Glaze is closed-source software whose perturbation budget cannot be exactly set to $16/255$, we set the intensity to High, the highest setting..
Transformations and Purifications. We evaluate the robustness of Dormant against 5 popular image transformations: JPEG compression, Gaussian blur, Gaussian noise, Median blur, and Bit squeeze. Additionally, we further conduct experiments on 6 advanced purification methods specifically designed to purify the added protective perturbation, including Impress [10], DiffPure [47], DDSPure [12], GrIDPure [83], Diffshortcut [43] and Noisy Upscaling [28].
Metrics. We evaluate the quality of generated videos using 6 image- and video-wise generative metrics used in prior human image animation methods [15, 68]. For the assessment of image-level quality, we report frame-wise LPIPS [79], Fréchet Inception Distance (FID) [25], Peak Signal to Noise Ratio (PSNR) [29] and Structural Similarity Index Measure (SSIM) [70]; while for video-level evaluation, we concatenate every consecutive 16 frames to form a sample and report FID-VID [6] and Fréchet Video Distance (FVD) [64], respectively.
@@ 4.2 Protection Performance
Overall Results. We evaluate the protection effectiveness and transferability of Dormant against eight pose-driven human image animation methods on the TikTok dataset, with the quantitative results shown in Table 1. Evaluation across six metrics demonstrates significant degradation in the quality of generated videos due to the protective effect of Dormant on all eight animation methods. Specifically, the average FID-VID, FVD, LPIPS, and FID increase from 48.03, 382.04, 0.30, and 65.54 to 148.45, 1212.88, 0.61 and 228.12, and the average PSNR and SSIM decrease from 17.01 and 0.731 to 13.22 and 0.418, respectively. Figure 3 visualizes the generated video frames revealing awful quality, including mismatches in the appearance of human identity, distorted backgrounds, and visually noticeable artifacts, etc. More visualization results for videos of Dormant against various animation techniques can be found in Figure 15 in Appendix A.2, which exhibit inconsistencies among frames. These results highlight the effectiveness of Dormant in preventing unauthorized image usage in human video generation, thereby safeguarding the individuals’ rights to portrait and privacy.
We also compare Dormant’s protection performance with six baseline methods that defend against unauthorized image and video generation. Quantitative results in Table 1 and qualitative results in Figure 3 demonstrate the superior protection performance of Dormant, showing the greatest effectiveness in degrading the quality of generated videos and the strongest transferability across different animation techniques. In contrast, the concurrent work VGMShield only provides acceptable results on the SVD-based method ControlNeXt while showing limited transferability to other animation methods. SDS has better PSNR on ControlNeXt and PhotoGuard has better FID-VID on MagicPose, possibly due to the resulting color changes in the generated videos. Glaze also produces visual artifacts in generated videos and has better FVD on Champ and UniAnimate. However, the facial features remain clear, leaving the individual’s portrait still vulnerable. Moreover, although these baseline protections may perform better on some metrics for certain animation methods, Dormant still 1) shows a clear advantage in other metrics for the corresponding animation method and 2) outperforms them across all other animation methods.
@@
Table 2: Quantitative results on various datasets.
@@
Results on Various Datasets. Besides the TikTok dataset, we also evaluate the protection performance of Dormant on three other datasets using Animate Anyone, including human dance generation on the Champ dataset, fashion video synthesis on the UBC Fashion dataset, and speech video generation on the TED Talks dataset. Quantitative results in Table 2 demonstrate the remarkable performance of Dormant in defending against pose-driven human image animation, with the average FID-VID, FVD, LPIPS and FID increasing from 53.78, 406.03, 0.22 and 60.81 to 166.28, 1288.20, 0.57 and 260.53, and the average PSNR and SSIM decreasing from 19.44 and 0.745 to 14.69 and 0.486, respectively. The poor-quality video frames shown in Figure 4 also validate the protection effectiveness of Dormant. More visualization results for videos can be found in Figure 16 in Appendix A.2.
@@
Results of Human and GPT-4o Studies. Besides using quantitative metrics for assessment, we also evaluate the performance of Dormant compared to baseline methods from the perspectives of human perception and multimodal large language model, GPT-4o [48]. In the human study, participants are provided with a reference video and five generated video options in GIF format. They are asked to select the one that best protects portrait and privacy rights compared to the reference. The five options include videos generated from images protected by Dormant and four relatively well-performing baseline methods (SDS, Glaze, VGMShield, and PhotoGuard) according to the results in Table 1. Our questionnaire consists of ten questions, and we ultimately collect 82 valid responses. As shown in Table 3, 74.26% of participants pick Dormant as the most effective protection method, highlighting its superiority from the perspective of human perception. Additionally, we also ask GPT-4o to rank the five frame sequences sampled from the generated videos, where Dormant achieves the highest average ranking of 1.5. Specific examples of human and GPT-4o studies can be found in Figure 13 and Figure 14 in Appendix A.1.
4.3 Protection Robustness
Figure 5: Qualitative results of transformations that most degrade the protection, with JPEG compression (quality=75), Gaussian blur (sigma=3), Gaussian noise (scale=0.05), Median blur (kernel size=9), and Bit squeeze (depth=3).
@@ 4.3 Protection Robustness
Robustness against Transformations. We apply five transformations with different parameter settings to both the original and protected images, to investigate their impact on the quality of generated videos and the effectiveness of protection. As shown in Figure 6, the results indicate that: 1) Transformations can only reduce the protection effect to some extent but cannot completely remove the impact of the protective perturbation. 2) Transformations also affect the quality of generated videos. In particular, increasing the transformation parameters to minimize the protection effect would also noticeably degrade the quality of the generated videos. Figure 5 presents the qualitative results when various transformations most degrade the protection effect. Even in these cases, the generated videos still exhibit low quality due to residual protective perturbation and the impact of excessive transformations, displaying issues including distorted backgrounds, mismatched faces and visual artifacts. These results validate the robustness of Dormant against various transformations.
@@
Robustness against Purifications. We evaluate the robustness of Dormant against six purification methods, and the results are shown in Table 4. While the protective effect of Dormant is somewhat reduced by these purifications, there still remains a noticeable quality gap between the resulting videos and those generated from the original images. Qualitative results are presented in Figure 7. Impress shows almost no purifying effect. DiffPure and DDSPure would introduce external strange patterns and light streaks into the generated videos. GrIDPure and Noisy Upscaling fail to restore detailed facial features. In contrast, while Diffshortcut utilizes CodeFormer[85] to restore human faces (which are vivid but differ from the reference), there are still noticeable artifacts in other regions. As a result, videos generated using protected images after purification cannot be used directly for malicious purposes due to their low quality, and the individual’s privacy and portrait rights are still effectively safeguarded.
@@
4.4 Ablation Study
Figure 8: Ablation study on the proposed objectives.
Impact of Perturbation Budget $\eta$ and PGD Iterations $N$. We conduct experiments by varying $\eta$ from $2/255$ to $32/255$ and $N$ from $50$ to $500$, to investigate their impact on protection performance. The results are presented in Figure 9. As the perturbation budget $\eta$ increases, the protection effect of Dormant becomes progressively stronger, and $\eta=8/255$ can already significantly degrade the quality of generated videos, resulting in FVD and FID values of 1158.46 and 240.83, respectively. This provides users with the flexibility to balance invisibility and effectiveness according to their needs. The increase in PGD iterations $N$ initially causes the protection effect to decrease and then increase, possibly due to overfitting, which leads the optimization to fall into poor local maxima and then escape. Optimizing for 50 iterations is sufficient to significantly degrade the quality of generated videos, with FVD and FID being 1522.35 and 272.61, thereby further reducing the time cost required for protection.
(a) Perturbation budget $\eta$
(b) PGD iterations $N$
Figure 9: Impact of budget $\eta$ and iterations $N$.
5 Discussion
5.1 Performance on Image-to-Video/Image
Besides pose-driven human image animation, we conduct additional experiments on image-to-video methods that do not rely on external guidance (SVD) and those that use text conditions for guidance (PIA, I2VGen-XL, and AnimateDiff). We also consider the scenario of image manipulation and evaluate the performance of Dormant using four image-to-image methods that use textual prompts as conditions for image editing. We utilize BLIP-2 [40] to generate captions corresponding to the reference images, resulting in prompts such as “A woman in a white dress”. To guide video generation, we add the suffix “is dancing” to these captions. For image-to-image methods, we use the original captions for SDEdit and InstructPix2Pix, while for LEDITS++ and DiffEdit, we generate variant prompts using GPT-4o, such as “A woman in a blue dress”. Since FID-VID and FVD cannot be calculated here, we employ two other metrics, CLIP-I and DINO, which compute the average pairwise cosine similarity between CLIP image embeddings [53] and ViT-S/16 DINO embeddings [14] of generated and reference images, respectively.
Table 5: Quantitative results on various resolutions of images.
[…]
256 $\bm{\times}$ 256 Clean 56.23 462.36 16.04 0.625 0.304 94.07
Protect 260.26 2165.94 13.19 0.306 0.596 285.03
256 $\bm{\times}$ 512 Clean 69.54 480.56 15.74 0.637 0.349 83.85
Protect 289.11 1769.24 12.06 0.322 0.640 292.09
512 $\bm{\times}$ 512 Clean 41.61 362.75 17.76 0.741 0.282 65.00
Protect 162.87 1364.00 11.82 0.374 0.639 245.06
512 $\bm{\times}$ 768 Clean 58.86 505.83 16.84 0.722 0.328 85.67
Protect 204.48 1419.64 11.62 0.403 0.653 243.46
768 $\bm{\times}$ 768 Clean 42.93 425.98 17.89 0.779 0.284 76.45
Protect 163.07 1423.93 10.58 0.441 0.650 189.82
@@ 5.1 Performance on Image-to-Video/Image
5.2 Performance on Commercial Services
We further evaluate the performance of Dormant on six real-world commercial services. To effectively defend against these closed-source generative models in a black-box manner, we set the budget $\eta$ to $32/255$ to generate protective perturbation. The results are presented in Table 7. Dormant significantly degrades the quality of generated videos and images, with the average CLIP-I, DINO, and SSIM decreasing from 0.804, 0.788, and 0.630 to 0.662, 0.521, and 0.475, and the average LPIPS and FID increasing from 0.474 and 145.22 to 0.643 and 269.64, respectively. As shown in Figure 12, generated videos and images suffer from poor quality, such as mismatched human appearance (Gen-2, DreaMoving, and Ying) and distorted visuals (Viva, NovelAI, and Scenario). These results validate that Dormant effectively safeguards portrait and privacy rights, showcasing its capability for real-world applications in preventing human image misuse.

### v2

@@ 3.4 Dormant
Overall, to optimize the protective perturbation $\delta$ for preventing unauthorized image usage in pose-driven human image animation, we propose $\mathcal{L}_{vae}$ and $\mathcal{L}_{feature}$ for feature misextraction, $\mathcal{L}_{frame}$ for frame incoherence. We also apply several techniques to further enhance the perturbation $\delta$ during optimization, including LPIPS [91] for imperceptibility, EoT [5] for robustness, and Momentum [19] for transferability.
@@
EoT. To enhance the robustness of $\delta$ against various transformations, we adopt EoT into the PGD process. EoT introduces a distribution of transformation functions to the input and optimizes the expectation of the objective function values over these transformations. To reduce time and computational costs, during each PGD step, we randomly sample one of the following transformations and apply it to the protected image $x_{p}$: Gaussian blur, JPEG compression, Gaussian noise, Random resize (resizing to a random resolution and then back to the original size), or no transformation. These selected transformations are commonly adopted as potential countermeasures in prior studies and have been shown to weaken the effect of protective perturbations [95, 3, 31].
@@ 4.1 Experimental Setup
Pose-driven Human Image Animation Methods. We comprehensively evaluate the effectiveness and transferability of Dormant on 8 cutting-edge and widely-used human image animation methods, including Animate Anyone [34], MagicAnimate [83], MagicPose [14], MusePose [74], Champ [98], MuseV [81], UniAnimate [79], and ControlNeXt [61]. These 8 methods represent SOTA open-source animation techniques within the last two years and cover all major types. Despite advances in controllability and realism, some methods have also been implicated in generating dancing deepfakes on video platforms, raising societal concerns [60, 49, 16].
Datasets. We conduct experiments on 4 datasets: TikTok [38], Champ [98], UBC Fashion [88], and TED Talks [73]. For the TikTok dataset, we follow the settings used in prior human image animation works [14, 78] and utilize their 10 TikTok-style videos showing different people from the web for evaluation. These videos contain between 248 and 690 frames. Specifically, we use the first frame of each video as the reference image and extract pose sequences from the remaining frames using DWPose [85] or DensePose [26] to guide the video generation. Additionally, we also sample 10 videos from the Champ training set, and the UBC Fashion and TED Talks test sets for evaluation, resulting in videos with 237-848, 303-355, and 135-259 frames, respectively. All video frames are resized to $512\times 512$, and we further conduct experiments on other image resolutions, random reference images, and jump-cut pose sequences in Appendix B.
Baseline Protections. We compare the protection performance of Dormant with 6 baseline methods defending against text/image-to-image generation (SDS [84], Glaze 2.1 [71], Mistv2 [96], PhotoGuard [70], and AntiDB [76]) and image-to-video generation (VGMShield [58]). For a fair comparison, we set the perturbation budget $\eta=16/255$ and PGD iterations $N=200$ for all protection methods [^4] Glaze is closed-source software whose perturbation budget cannot be exactly set to $16/255$, we set the intensity to High, the highest setting.. We evaluate image similarity before and after protections in Appendix G.
Transformations and Purifications. We evaluate the robustness of Dormant against 5 popular image transformations: JPEG compression, Gaussian blur, Gaussian noise, Median blur, and Bit squeeze. Additionally, we further conduct experiments on 6 advanced purification methods specifically designed to purify the added protective perturbation, including Impress [10], DiffPure [55], DDSPure [11], GrIDPure [95], Diffshortcut [51], and Noisy Upscaling [31].
Metrics. We evaluate the quality of generated videos using 6 image- and video-wise generative metrics used in prior human image animation methods [14, 78]. For the assessment of image-level quality, we report frame-wise LPIPS [91], Fréchet Inception Distance (FID) [28], Peak Signal to Noise Ratio (PSNR) [32], and Structural Similarity Index Measure (SSIM) [80]; while for video-level evaluation, we concatenate every consecutive 16 frames to form a sample and report FID-VID [6] and Fréchet Video Distance (FVD) [75], respectively.
@@
We provide detailed introductions to the evaluated animations, protections, transformations, purifications, and image-to-video and image-to-image methods in Appendix A.
@@ 4.2 Protection Performance
Overall Results. We evaluate the protection effectiveness and transferability of Dormant against eight pose-driven human image animation methods on the TikTok dataset, with the quantitative results shown in Table 1. Evaluation across six metrics demonstrates significant degradation in the quality of generated videos due to the protective effect of Dormant on all eight animation methods. Specifically, the average FID-VID, FVD, LPIPS, and FID increase from 48.03, 382.04, 0.30, and 65.54 to 148.45, 1212.88, 0.61 and 228.12, and the average PSNR and SSIM decrease from 17.01 and 0.731 to 13.22 and 0.418, respectively. We show non-cherry-picked generated video frames in Figure 3, which display mismatched identities, distorted backgrounds, and visual artifacts. More visualization results for videos of Dormant against various animation techniques can be found in Figure 20 in Appendix H, which exhibit inconsistencies among frames. These results highlight the effectiveness of Dormant in preventing unauthorized image usage in human video generation, thereby safeguarding individuals’ rights to portrait and privacy.
We also compare Dormant’s protection performance with six baseline methods that defend against unauthorized image and video generation. Quantitative results in Table 1 and qualitative results in Figure 3 demonstrate the superior protection performance of Dormant, showing the greatest effectiveness in degrading the quality of generated videos. While Dormant exhibits the strongest transferability and performs effectively across all 8 animation techniques, baseline protections show limited transferability and are only effective for certain animation methods, yielding some better metrics with different protective effects. Specifically, the concurrent work VGMShield achieves slightly higher LPIPS on the SVD-based method ControlNeXt but lacks transferability to other animation methods. SDS tends to lighten colors and smooth out details, resulting in better PSNR on ControlNeXt. PhotoGuard induces noticeable inconsistencies between adjacent frames, showing slightly higher FID-VID on MagicPose. Glaze produces visual artifacts and intensifies colors in generated videos, leading to better FVD on Champ and UniAnimate. Moreover, although these baseline protections may perform better on some metrics for certain animation methods, Dormant still 1) shows a clear advantage in other metrics for the corresponding animation method; and 2) outperforms them across all other animation methods.
@@
Results on Various Datasets. Besides the TikTok dataset, we also evaluate the protection performance of Dormant on three other datasets using Animate Anyone, including human dance generation on the Champ dataset, fashion video synthesis on the UBC Fashion dataset, and speech video generation on the TED Talks dataset. Quantitative results in Table 2 demonstrate the remarkable performance of Dormant in defending against pose-driven human image animation, with the average FID-VID, FVD, LPIPS and FID increasing from 53.78, 406.03, 0.22 and 60.81 to 166.28, 1288.20, 0.57 and 260.53, and the average PSNR and SSIM decreasing from 19.44 and 0.745 to 14.69 and 0.486, respectively. The poor-quality video frames shown in Figure 4 also validate the protection effectiveness of Dormant. More visualization results for videos can be found in Figure 21 in Appendix H.
4.3 Human and GPT-4o Studies
Human Study. To bridge the gap between quantitative metrics and human perception, as well as study whether Dormant outperforms baseline methods in protecting privacy and portrait rights from the perspective of real-world audiences, we conduct a survey study. Our questionnaire consists of two parts: 1) demographic data collection, including five questions on age, gender, education, expertise, and familiarity; and 2) ten questions asking participants to select the video that best protects portrait and privacy rights compared to the reference, each question containing a reference video and five videos generated from images protected by Dormant and four baseline methods (SDS, Glaze, VGMShield, and PhotoGuard). The videos used are the 10 videos in the TikTok test set, derived from experiments conducted using Animate Anyone, with the corresponding metrics provided in Table 1. The detailed questionnaire can be found in Appendix F.1.
Table 2: Protection performance on various datasets.
@@ 4.3 Human and GPT-4o Studies
We distributed the questionnaire online and recruited participants via social media, offering lottery-based compensation upon completion. A total of 82 valid responses were collected. The participants’ ages primarily ranged from 18 to 54, with the majority (63.41%) falling between 25 and 34. Among the participants, 57.32% were male and 40.24% were female; 80.48% held or were pursuing a bachelor’s degree or higher; 39.02% had a background in computer-related fields; and 70.73% had at least heard of image-to-video generation or were more familiar with it. As shown in Table 3, Dormant achieved an overall pick rate of 74.27% (609/820), significantly outperforming baseline protections. Detailed results for each question are shown in Figure 15(a) in Appendix F.3. Notably, 67 participants selected Dormant for at least five questions, including 29 with a computer-related background (32 in total). 37 participants selected Dormant for all ten questions, while only 8 participants did not choose Dormant for any of the ten questions, with SDS being their preferred method, yielding a pick rate of 72.50% (58/80). Of these eight participants, only one had heard of image-to-video generation, and the others were less familiar with it.
@@
GPT-4o Study. As LLM-as-a-Judge is now widely adopted in many complex tasks [25], we also use GPT-4o [56] to simulate the above human study. GPT-4o is a multimodal large language model capable of understanding image semantics and perceiving pixel-level differences. We input GPT-4o with a reference and five generated frame sequences, asking it to rank them based on their effectiveness in protecting portrait and privacy rights, focusing mainly on factors such as appearance mismatches and background changes that differentiate the generated frames from the reference. GPT-4o is instructed to analyze step by step and output both the ranking and its reasoning. Our prompt is detailed in Appendix F.2. We further manually verify all outputs from GPT-4o to ensure there are no hallucinations. As shown in Table 3, Dormant achieves the highest average ranking of 1.5, and is consistently ranked first or second across all samples (see Figure 15(b) in Appendix F.3). While GPT-4o ranks Dormant first, it notes that this method “makes the individual unrecognizable and drastically alters the background”. Glaze is preferred by GPT-4o when Dormant is ranked second, as it “incorporates more prominent and deliberate texture patterns”.
Figure 6: Qualitative results of transformations that most degrade the protection, with JPEG compression (quality=75), Gaussian blur (sigma=3), Gaussian noise (scale=0.05), Median blur (kernel size=9) and Bit squeeze (depth=3).
4.4 Protection Robustness
Robustness against Transformations. We evaluate five transformations commonly used in prior studies [95, 31, 18], including three transformations applied in EoT (JPEG compression, Gaussian blur, and Gaussian noise) and two unseen transformations (Median blur and Bit squeeze). We apply them with different parameter settings to both the original and protected images to investigate their impact on the quality of generated videos and the effectiveness of protection. As shown in Figure 5 (with the red dotted line representing the results of videos generated using the original image without transformations), the results indicate that: 1) transformations can only reduce the protection effect to some extent but cannot completely remove the impact of the protective perturbation; 2) transformations also affect the quality of generated videos. In particular, increasing the transformation parameters to minimize the protection effect would also noticeably degrade the quality of the generated videos. Figure 6 presents the qualitative results when various transformations most degrade the protection effect. Even in these cases, the generated videos still exhibit low quality due to residual protective perturbation and the impact of excessive transformations, displaying issues including distorted backgrounds, mismatched faces, and visual artifacts. These results validate the robustness of Dormant against various transformations.
@@ 4.4 Protection Robustness
Robustness against Purifications. We evaluate the robustness of Dormant against six purification methods, and the results are shown in Table 4. While the protective effect of Dormant is somewhat reduced by these purifications, there still remains a noticeable quality gap between the resulting videos and those generated from the original images. Qualitative results are presented in Figure 7. Impress shows almost no purifying effect. DiffPure and DDSPure would introduce external strange patterns and light streaks into the generated videos. GrIDPure and Noisy Upscaling fail to restore detailed facial features. In contrast, while Diffshortcut utilizes CodeFormer [97] to restore human faces (which are vivid but differ from the reference), there are still noticeable artifacts in other regions. As a result, videos generated using protected images after purification cannot be used directly for malicious purposes due to their low quality, and the individuals’ privacy and portrait rights are still effectively safeguarded.
@@
4.5 Protection Transferability
Results on Image-to-Video/Image. Besides human image animation methods that use pose sequences to guide video generation from the reference image, we conduct additional experiments on four other types of image-to-video methods that directly generate videos from the given image without external guidance (SVD) and those that use text prompts for guidance (PIA, I2VGen-XL, and AnimateDiff). Moreover, we also consider the scenario of image manipulation and evaluate the performance of Dormant using four image-to-image methods that use text prompts as conditions for image editing. We utilize BLIP-2 [44] to generate captions corresponding to the reference images, resulting in prompts such as “A woman in a white dress”. To guide video generation, we add the suffix “is dancing” to these captions. For image-to-image methods, we use the original captions for SDEdit and InstructPix2Pix, while for LEDITS++ and DiffEdit, we generate variant prompts using GPT-4o, such as “A woman in a blue dress”. Since FID-VID and FVD cannot be calculated here, we employ two other metrics, CLIP-I and DINO, which compute the average pairwise cosine similarity between CLIP image embeddings [62] and ViT-S/16 DINO embeddings [13] of generated and reference images, respectively.
Figure 10: Qualitative results on commercial services.
@@ 4.5 Protection Transferability
Results on Commercial Services. We further evaluate the performance of Dormant on six real-world commercial services. To effectively defend against these closed-source generative models in a black-box manner, we set the budget $\eta$ to $32/255$ to generate protective perturbation. The results are presented in Table 6. Dormant significantly degrades the quality of generated videos and images, with the average CLIP-I, DINO, and SSIM decreasing from 0.804, 0.788, and 0.630 to 0.662, 0.521, and 0.475, and the average LPIPS and FID increasing from 0.474 and 145.22 to 0.643 and 269.64, respectively. As shown in Figure 10, generated videos and images suffer from poor quality, such as mismatched human appearance (Gen-2, DreaMoving, and Ying) and distorted visuals (Viva, NovelAI, and Scenario). These results validate that Dormant effectively safeguards portrait and privacy rights, showcasing its capability for real-world applications in preventing human image misuse.
4.6 Ablation Study
Figure 11: Ablation study on the proposed objectives.
We further study the impact of using other feature extractors in $\mathcal{L}_{feature}$. As mentioned in Section 3.2, we use the CLIP image encoder to extract semantic features and ReferenceNets to capture fine-grained details from the image. We replace our design of CLIP + ReferenceNets with DINO [13] + BLIP image encoder [45], yielding FVD and FID values of 1080.09 and 153.55 for the generated videos. This result outperforms most baseline protections (except Glaze) but falls short of Dormant by a noticeable margin, further highlighting the importance of inducing misextraction of appearance features and also the superiority of our design for feature extractors.
Impact of Perturbation Budget $\eta$ and PGD Iterations $N$. We conduct experiments by varying $\eta$ from $2/255$ to $32/255$ and $N$ from $50$ to $500$, to investigate their impact on protection performance. The results are presented in Figure 12. As the perturbation budget $\eta$ increases, the protection effect of Dormant becomes progressively stronger, and $\eta=8/255$ can already significantly degrade the quality of generated videos, resulting in FVD and FID values of 1158.46 and 240.83, respectively. This provides users with the flexibility to balance invisibility and effectiveness according to their needs. The increase in PGD iterations $N$ initially causes the protection effect to decrease and then increase, possibly due to overfitting, which leads the optimization to fall into poor local maxima and then escape. Optimizing for 50 iterations is sufficient to significantly degrade the quality of generated videos, with FVD and FID being 1522.35 and 272.61, thereby further reducing the time cost required for protection. Impact of pose repeats $F$ and decay factor $\mu$ is provided in Appendix D.
(a) Perturbation budget $\eta$
(b) PGD iterations $N$
Figure 12: Impact of budget $\eta$ and iterations $N$.
5 Discussion
Adaptive Attack. We propose an adaptive attack under the assumption that the attacker has access to multiple protected images of the victim. While animation methods typically take a single reference image as input to generate videos, the attacker can utilize these images to craft a refined reference. Inspired by [59] which reveals Glaze’s vulnerability to linear interpolation and pixel averaging, we design a strategy to create a purified image from multiple protected images. Specifically, we apply linear interpolation to adjacent pairs of five protected images spaced five frames apart, followed by pixel averaging of the four interpolated results to produce the final image for video generation, with the resulting FID-FVD, FVD, LPIPS, FID, PSNR, and SSIM on the TikTok dataset being 69.18, 728.89, 0.474, 162.50, 17.06, and 0.641, respectively. While the protective effect is partially reduced, the generated videos remain of noticeably low quality. More adaptive attacks are provided in Appendix C.
Application Scenarios. Dormant applies protective perturbations to human images to prevent potential unauthorized usage in pose-driven animation methods. It is primarily targeted at users who: 1) may place human images in uncontrollable environments like social media or the web; and 2) prioritize privacy and portrait rights and accept sacrificing a bit of image quality for protection. Dormant may not be directly suitable for legitimate use cases that involve processing or editing original images for valid purposes, such as photo retouching or authorized content creation.
Anomalous Cases. We examine cases where baseline protections outperform Dormant across at least four animations on the TikTok dataset. While Dormant exhibits sufficient effectiveness, these methods yield even better protective effects on certain samples. Specifically, Glaze performs better on sample 002, producing highly exaggerated colors and visible speckles in the generated videos. PhotoGuard and SDS perform better on sample 006 which features sharply contrasting colors, by lightening the colors and blurring sharpness and details in the generated videos. We present the FVD for experiments on Animate Anyone and MagicPose as examples in Figure 13, and qualitative results in Figure 15 in Appendix E.
(a) Results of Animate Anyone
(b) Results of MagicPose
Figure 13: Results of experiments on Animate Anyone and MagicPose, with red boxes marking samples where baseline protections perform better across both animation methods.
We also examine failure cases where Dormant performs relatively poorly against the commercial service Gen-2 and the latest Gen-3. As shown in Figure 16 in Appendix E, Dormant mainly induces identity mismatch but introduces minimal distortion, and the generated videos remain realistic. Image-to-video models applied in commercial services are typically trained on larger, higher-quality datasets with more parameters. Exploring how to get feedback from these advanced models via black-box access to improve Dormant’s transferability would be meaningful for future work.
Limitations and Future Works. A key limitation of Dormant is that its effectiveness is validated based on empirical evaluation, rather than being provably secure. Additionally, Dormant faces the challenge of being future-proof, as image-to-video generation is continuously evolving. While Dormant has shown effectiveness and is expected to remain effective against existing feature extractors used in current open-source SOTA animation methods that are driven by pose, its performance against future paradigms in feature extraction and different video generation techniques guided by other modal conditions requires ongoing validation and improvements. Future work could explore the use of expert models to specifically extract disentangled image features (e.g., human face, body posture, and background) to induce feature misextraction, and integrate guidance from additional modalities to create frame incoherence during optimization.